### v1

# Standards-Compliant DM-RS Allocation via Temporal Channel Prediction for Massive MIMO Systems

Abstract
Reducing feedback overhead in beyond 5G networks is a critical challenge, as the growing number of antennas in modern massive MIMO systems substantially increases the channel state information (CSI) feedback demand in frequency division duplex (FDD) systems. To address this, extensive research has focused on CSI compression and prediction, with neural network-based approaches gaining momentum and being considered for integration into the 3GPP 5G-Advanced standards. While deep learning has been effectively applied to CSI-limited beamforming and handover optimization, reference signal allocation under such constraints remains surprisingly underexplored. To fill this gap, we introduce the concept of channel prediction-based reference signal allocation (CPRS), which jointly optimizes channel prediction and DM-RS allocation to improve data throughput without requiring CSI feedback. We further propose a standards-compliant ViViT/CNN-based architecture that implements CPRS by treating evolving CSI matrices as sequential image-like data, enabling efficient and adaptive transmission in dynamic environments. Simulation results using ray-tracing channel data generated in NVIDIA Sionna validate the proposed method, showing up to 36.60% throughput improvement over benchmark strategies.
Index Terms:
5G NR, Massive MIMO, Reference Signal, DM-RS, Channel Prediction, Deep Learning.
I Introduction
The multiple-input-multiple-output (MIMO) systems [GESBERT2002] have been at the core of innovation since 4G LTE to enhance data throughput. However, despite its original goal, the introduction of massive MIMO has led to the challenge of increased channel dimensionality [LU2014]. As the dimension of the channel matrix increases, both the transmission of the reference signal for channel estimation and the feedback of downlink channel state information (CSI) to the base station have become significant sources of overhead. The downlink CSI is critical for enhancing communication tasks such as beamforming [JAFAR2001], handover [TOLLI2008], and reference signal allocation [CAI2005]. In time division duplexing (TDD) systems, the downlink channel can be inferred from the uplink channel due to channel reciprocity. However, in frequency division duplexing (FDD) systems, CSI feedback is required [LOVE2008]. As CSI feedback has become a major source of overhead in FDD systems, extensive research has been conducted to address this challenge [SHEN2017].
Figure 1: The NR frame structure and its corresponding transmission timings are defined based on a 10ms radio frame, consisting of ten 1ms subframes. The number of slots per subframe, N, is flexibly determined by the numerology specified in 3GPP TS 38.211 table 4.3.2-1 [TS138211], where $\textnormal{N}=2^{\mu}$, $\mu\in\mathbb{Z}$, and $0\leq\mu\leq 6$. Within each slot, the PDSCH DM-RS can be configured to occupy 1 to 4 out of the 14 OFDM symbols.
Recently, deep learning has emerged as a breakthrough solution for reducing CSI feedback overhead and is also being discussed in the 5G Advanced standards following 3GPP Rel-18 [RP213599]. The solutions can be broadly classified into two categories: CSI compression [WEN2018] and channel prediction [WANGJIE2019]. We focus on channel prediction, which estimates downlink CSI from uplink CSI or data signals, even without explicit CSI feedback. The network architectures have evolved from CNNs [WANGJIE2019] to LSTMs [HELMY2023], and more recently, transformers [ZHOU2024]. There have also been attempts to jointly optimize beamforming and handover, the primary objectives of channel prediction. For instance, [CHU2022] proposed a deep reinforcement learning algorithm for channel prediction and beamforming, while [JIANG2022] used a transformer-based approach for beamforming with predicted CSI. In terms of handover, channel quality prediction has been used to guide handover decisions [SKABA2024]. However, to the best of the authors’ knowledge, no prior study has explored the critical issue of reference signal allocation in conjunction with channel prediction, despite their strong interdependence.
We address the allocation of the physical downlink shared channel (PDSCH) demodulation reference signal (DM-RS) within the NR frame structure, as illustrated in Fig. 1. Among various reference signals defined in 3GPP NR, the PDSCH DM-RS plays a critical role in determining data throughput, as it directly enables the user equipment (UE) to estimate the channel for data demodulation. Moreover, unlike other reference signals transmitted over the control plane, the DM-RS is carried on the data plane alongside data symbols. The NR standard provides various configuration options for the flexible placement of 1 to 4 DM-RS symbols within each slot, the transmission time interval (TTI) of the NR frame, as detailed in Table I, offering a level of flexibility not available in previous LTE standards. This design accommodates practical challenges such as mobility-induced channel aging, noise, and estimation uncertainty. The in-slot placement makes DM-RS allocation particularly important, as increasing the number of DM-RS symbols improves channel estimation accuracy at the expense of reducing transmittable data symbols, highlighting an inherent trade-off in the allocation design. Therefore, the DM-RS must be optimally allocated to balance channel estimation quality and the number of available data symbols, considering the current downlink CSI. However, as the standard does not specify which allocation to use under specific downlink CSI conditions, network operators have the flexibility to configure it based on their own channel assessments.
TABLE I: DM-RS positions within a 14-symbol slot duration are specified in Table 7.4.1.1.2-3/4 of 3GPP TS 38.211 [TS138211], with $l_{0}\in\{2,3\}$ and $l_{1}\in\{11,12\}$. Only Type-A configurations are included, as Type-B applies to a special use case of mini-slot-based scheduling with fewer supported allocation patterns.
DM-RS Length DM-RS positions pos0 pos1 pos2 pos3 Single-Symbol $l_{0}$ $l_{0}$, $l_{1}$ $l_{0}$, 7, 11 $l_{0}$, 5, 8, 11 Double-Symbol $l_{0}$ $l_{0}$, 10 - -
This paper proposes a channel prediction-based reference signal allocation method, with a focus on DM-RS, for massive MIMO systems in B5G networks. While designed with B5G scenarios in mind, the proposed method remains fully compatible with the 5G NR frame structure and is applicable under current standard specifications.
Contribution: 1) This paper introduces the concept of channel prediction-based reference signal allocation (CPRS), with a particular focus on DM-RS, as a practical approach to reducing CSI feedback overhead in massive MIMO FDD scenarios. 2) We propose a ViViT/CNN-based [ARNAB2021] CPRS algorithm, designed with consideration of the NR frame structure, which interprets time-varying uplink channel matrices as video data. 3) We generate ray-tracing-based channel data and perform simulations using NVIDIA Sionna [NVIDIA], demonstrating the effectiveness of the proposed method.
II System Model and Problem Formulation
We consider a massive MIMO-OFDM communication system, focusing on a downlink scenario with UE mobility, where the BS and UE communicate based on the NR frame structure. The numbers of receive antennas, transmit antennas, and subcarriers are denoted by $N_{\textnormal{R}}$, $N_{\textnormal{T}}$, and $N_{\textnormal{C}}$, respectively. The antennas at the UE receive the same transmitted signal, and the received signal at the UE for the $t$-th OFDM symbol in the slot and $k$-th subcarrier can be expressed as follows:
$y_{t,k}=\mathbf{h}_{\textnormal{DL},t,k}\,\mathbf{v}_{t,k}\,x_{t,k}+\mathbf{w}_{t,k},$ (1)
where $\mathbf{h}_{\textnormal{DL},t,k}\in\mathbb{C}^{N_{\textnormal{R}}\times N_{\textnormal{T}}}$ is the downlink channel matrix associated with the $t$-th symbol and $k$-th subcarrier, $\mathbf{v}_{t,k}\in\mathbb{C}^{N_{\textnormal{T}}\times 1}$ is the precoding vector, $x_{t,k}\in\mathbb{C}$ is the transmit symbol, and $\mathbf{w}_{t,k}\in\mathbb{C}^{N_{\textnormal{R}}\times 1}$ denotes additive noise. Then, each $t$-th symbol within the slot has a distinct downlink CSI as follows:
$\mathbf{H}_{\textnormal{DL},t}=\left[\mathbf{h}_{\textnormal{DL},t,1}\,\mathbf{h}_{\textnormal{DL},t,2}\,\cdots\,\mathbf{h}_{\textnormal{DL},t,N_{{\textnormal{C}}}}\right]\in\mathbb{C}^{N_{\textnormal{R}}\times N_{\textnormal{T}}\times N_{\textnormal{C}}}$ (2)
We define the overall downlink CSI for a slot as $\mathbf{H}_{\textnormal{DL}}$, which consists of the CSI matrices corresponding to all 14 OFDM symbols in the slot. It is expressed as:
$\mathbf{H}_{\textnormal{DL}}=\left\{\mathbf{H}_{\textnormal{DL},t}\right\}_{t=1}^{14}$ (3)
In our system, the BS is assumed to predict $\mathbf{H}_{\textnormal{DL}}$ using the uplink CSI from the current slot, denoted as $\mathbf{H}_{\textnormal{UL}}=\{\mathbf{H}_{\textnormal{UL},i}\in\mathbb{C}^{N_{\textnormal{T}}\times N_{\textnormal{R}}\times N_{\textnormal{C}}}\}_{i\in I_{u}}$, where $I_{u}$ is the set of DM-RS symbol indices in the current uplink slot. We assume that the uplink DM-RS is transmitted at four symbols per slot and $I_{u}=\{2,5,8,11\}$. The downlink channel prediction using a predictor function $f(\cdot)$ is formulated as:
$\hat{\mathbf{H}}_{\textnormal{DL}}=f(\mathbf{H}_{\textnormal{UL}})$ (4)
The BS utilizes $\hat{\mathbf{H}}_{\textnormal{DL}}$ to identify the optimal allocation $p^{*}\in\mathcal{P}$, where $\mathcal{P}$ denotes the set of possible allocation configurations from which the BS selects one to configure the next slot, in order to maximize data throughput. We refer to this process as channel prediction-based reference signal allocation (CPRS), and define the corresponding optimization problem as follows:
$p^{*}=\operatorname*{arg\,max}_{p\in\mathcal{P}}\>\left\{R\cdot\textnormal{D}_{\textnormal{sent}}\cdot\left(1-\frac{n_{p}}{n_{s}}\right)\cdot\left(1-\textnormal{BLER}\right)\right\},$ (5)
Figure 2: The proposed ViViT-based CPRS network captures the temporal variation of uplink channel matrices $\mathbf{H}_{\textnormal{UL}}$, obtained from DM-RS symbols at positions $I_{u}=\{2,5,8,11\}$ within an uplink slot. Leveraging a transformer architecture optimized for video data, it predicts the optimal DM-RS allocation for the next slot.
given $\hat{\mathbf{H}}_{\textnormal{DL}}=f(\mathbf{H}_{\textnormal{UL}})$, where $R$ is the code rate, $\textnormal{D}_{\textnormal{sent}}$ is the total number of bits transmitted in one slot, $n_{p}$ is the number of DM-RS symbols in allocation $p$, $n_{s}=14$ is the number of symbols within a slot, and BLER denotes the block error rate at the UE. The objective function of (5) represents the data throughput (bits per slot), following the model in [DAS2015]. We focus on the core optimization problem, excluding retransmissions as they fall outside the scope of this study. Here, $R$, $\textnormal{D}_{\textnormal{sent}}$, and $n_{s}$ are constants for a fixed numerology and thus independent of $p$, while $n_{p}$, the number of DM-RS symbols, is directly determined by the allocation $p$. The key challenge lies in modeling the complex relationship among $p$, $\mathbf{H}{\textnormal{UL}}$, and the resulting BLER at the UE in the next slot.
Even if the BS accurately estimates the $\mathbf{H}_{\textnormal{DL}}$, it remains challenging to determine the extent of BLER caused by channel errors in symbols between DM-RS placements. We adopt a data-driven approach to determine the relationship among $p$, $\mathbf{H}_{\textnormal{UL}}$, BLER, and ultimately, data throughput. Utilizing neural network algorithms often applied in nonconvex optimization, we propose an end-to-end method that jointly optimizes both the channel prediction problem in (4) and the DM-RS allocation in (5), classifying optimal allocation $p^{*}\in\mathcal{P}$ directly from $\mathbf{H}_{\textnormal{UL}}$. For the neural network-based classifier $\mathcal{F}(\cdot;\Theta)$ and estimated allocation $\hat{p}^{*}\in\mathcal{P}$, the proposed CPRS method is defined as:
$\hat{p}^{*}=\mathcal{F}(\mathbf{H}_{\textnormal{UL}};\Theta)$ (6)
We assume the possible allocation set $\mathcal{P}$ follows the NR standard in Table I. Without loss of generality, to simplify our analysis in this study, we consider the typical case of $l_{1}=11$, yielding $|\mathcal{P}|=13$ possible allocations, and apply Kronecker placement to focus on time-domain symbol allocation, as frequency-domain patterns (Type 1 and Type 2) are limited to only two fixed configurations and offer little flexibility.
III ViViT/CNN-based CPRS Network
Fig. 2 illustrates our proposed ViViT-based CPRS network, designed to predict optimal DM-RS allocations by exploiting the temporal evolution of uplink channel matrices $\mathbf{H}_{\textbf{UL}}$. We also introduce a lower-complexity CNN-based variant with the same input–output configuration, whose standard architecture is omitted for brevity.
Each $\mathbf{H}_{\textbf{UL},i}$ for $i\in\{2,5,8,11\}$ is first decomposed into real and imaginary parts, which are concatenated to form a multi-channel representation, then reshaped into non-overlapping patches that are flattened and embedded as tokens. Specifically, the channel matrix, originally of size $N_{\textnormal{T}}\times N_{\textnormal{R}}\times N_{\textnormal{C}}$, is reshaped by flattening the antenna dimensions, resulting in the following input tensor $\mathbf{X}_{i}$ for the network:
$\mathbf{H}_{\textbf{UL},i}\in\mathbb{C}^{N_{\textnormal{T}}\times N_{\textnormal{R}}\times N_{\textnormal{C}}}\quad\Longrightarrow\quad\mathbf{X}_{i}\in\mathbb{R}^{N_{\textnormal{T}}N_{\textnormal{R}}\times N_{\textnormal{C}}\times 2}$ (7)
This transformation enables the network to process spatial features jointly along the flattened antenna dimension, while preserving the temporal structure along the subcarrier axis $N_{\textnormal{C}}$. Moreover, it maintains compatibility with standard convolutional operations without losing the inherent spatial correlations present in the original matrix.
We then partition $\mathbf{X}_{i}$ into a set of non-overlapping patches and flatten each patch to form the initial tokens:
$\mathbf{z}_{i}^{(0)}=\bigl{\{}\mathbf{x}_{i,1},\,\mathbf{x}_{i,2},\,\dots,\mathbf{x}_{i,P}\bigr{\}},$ (8)
where $\mathbf{x}_{i,p}$ denotes the $p$-th patch and $P$ is the total number of patches. Learnable positional encodings are applied to the patch embeddings to incorporate spatial order.
Spatial Transformer Encoder
In the spatial transformer stage, each set of token embeddings $\mathbf{z}_{i}^{(0)}$ for a given time index $i$ passes through a standard transformer encoder block. Concretely, let $\mathbf{Z}_{i}^{(\ell)}\in\mathbb{R}^{P\times D}$ be the output of the $\ell$-th encoder layer, where $D$ is the embedding dimension of each token. Each encoder layer consists of multi-head self-attention (MHSA) and a position-wise feedforward network (FFN), each followed by layer normalization. The MHSA operation is defined by:
$\text{Attention}(Q,K,V)=\text{softmax}\!\Bigl{(}\frac{QK^{\top}}{\sqrt{d}}\Bigr{)}\!V,$ (9)
where $Q$, $K$, and $V$ are linear projections of the input $\mathbf{Z}_{i}^{(\ell-1)}$, and $d$ is the dimensionality of each head. Multiple attention heads are concatenated to capture diverse representation subspaces. After processing all four channel matrices in a slot independently in the spatial transformer, we obtain spatially enhanced tokens for each time $i$.
Temporal Transformer Encoder
We then stack the spatially transformed tokens across four time indices and feed them into the temporal transformer encoder. The same MHSA and FFN blocks now operate along the temporal dimension to learn correlations across DM-RS time positions. Stacking multiple temporal attention layers allows the network to capture channel evolution over time slots.
After the temporal transformer encoder, the final token representations are pooled and passed through a multi-layer perceptron (MLP) head to predict the optimal downlink DM-RS allocation index. This prediction leverages the spatio-temporal correlations learned via the transformer’s self-attention mechanism, which effectively captures long-range dependencies, to complete the end-to-end decision pipeline. The entire ViViT/CNN-based CPRS network is trained using a cross-entropy loss, aligning the model’s predictions with the ground-truth optimal DM-RS allocation labels.
The total computational complexity of the proposed ViViT model is $\mathcal{O}(L\cdot(N^{2}D+ND^{2}))$, where $L$ denotes the number of transformer layers, $N$ is the number of input tokens, and $D$ is the embedding dimension. For comparison, the CNN-based baseline consists of three 3D convolutional layers, each with computational complexity $\mathcal{O}(F\cdot C\cdot K^{3}\cdot H\cdot W\cdot T)$, where $K$ is the kernel size, $C$ the number of input channels, $F$ the number of output filters, and $H\times W\times T$ the spatial-temporal dimensions. Based on FLOP-level complexity analysis and efficient quantization techniques such as [ZMORA2021], both models are expected to be deployable at the microsecond level in a slot-based real-time system with parallel implementation.  [^1] With $L{=}8$, $N{=}64$, and $D{=}192$ as in our experiments, the ViViT model requires $\sim\!239.2\times 10^{6}$ FLOPs. Assuming INT8 quantization on an NVIDIA A100 (624 TFLOPs), the theoretical inference time is $\sim\!3.83~{}\mu\text{s}$. The CNN model, with fewer FLOPs, can run even faster.
III-A Dataset Generation
(a) Scene
(b) Ray Tracing
Figure 3: Process of obtaining channel matrices via ray tracing using NVIDIA Sionna [NVIDIA]. After constructing a dynamic scene that reflects movement as illustrated in (a), perform ray tracing by positioning the BS and UE as shown in (b).
We utilized the NVIDIA Sionna Ray Tracing to generate channel data [NVIDIA], as it more accurately reflects real-world propagation characteristics and physical environments than statistical channel models. Specifically, we employed the simple_street_canyon_with_cars scene included in the framework. As illustrated in Fig. 3, this scenario features a road between buildings, where vehicles move along both lanes. All vehicles move at a uniform speed $v\in\{60,65,\dots,80\}$ km/h. The BS is installed on one of the buildings and the UE is positioned in one of the vehicles. The BS has an 8×8 antenna array, while the UE has a 2×1 antenna array, forming a massive MIMO system. We consider an FDD scenario operating in NR band n1, as specified in Table 5.2-1 of [TS138101], using uplink and downlink carrier frequencies of 1.95 GHz and 2.14 GHz, respectively. The numerology is fixed to N=1, such that one subframe corresponds to a single slot. The maximum number of rays was set to 20, with a maximum ray depth of 4 and a bandwidth of 100 MHz.
For each velocity $v$, we obtained the uplink and downlink channel matrices for every symbol within 400 consecutive slots, each comprising 14 symbols and 64 subcarriers. Additionally, we augmented the original dataset by adding Gaussian noise with a magnitude of approximately 0.05 relative to the original channel matrices, in order to emulate practical channel estimation errors caused by uplink noise and other impairments. This augmentation generated an additional 4,000 slots. This resulted in a total dataset comprising 6,000 slots. Subsequently, for each slot, the optimal DM-RS allocation index was determined via exhaustive evaluation of all candidate configurations using Sionna link-level simulations, based on the objective function defined in (5). The dataset was then divided into training, validation, and test sets using an 8:1:1 ratio.
TABLE II: Classification accuracy comparison of different schemes for selecting the optimal DM-RS allocation index in the next slot, evaluated over 250 test slots. The results are reported per SNR value, excluding SNRs below –5 dB where all DM-RS allocations yield zero data throughput.

### v2

@@
 Abstract
-Reducing feedback overhead in beyond 5G networks is a critical challenge, as the growing number of antennas in modern massive MIMO systems substantially increases the channel state information (CSI) feedback demand in frequency division duplex (FDD) systems.
+Reducing CSI feedback overhead in beyond 5G networks is a critical challenge.
+The growing number of antennas in modern massive MIMO systems substantially increases the channel state information (CSI) feedback demand in frequency division duplex (FDD) systems.
 To address this, extensive research has focused on CSI compression and prediction, with neural network-based approaches gaining momentum and being considered for integration into the 3GPP 5G-Advanced standards.
-While deep learning has been effectively applied to CSI-limited beamforming and handover optimization, reference signal allocation under such constraints remains surprisingly underexplored.
+While deep learning has been effectively applied to CSI-limited beamforming and handover optimization, reference signal allocation under such constraints remains comparatively underexplored.
 To fill this gap, we introduce the concept of channel prediction-based reference signal allocation (CPRS), which jointly optimizes channel prediction and DM-RS allocation to improve data throughput without requiring CSI feedback.
-We further propose a standards-compliant ViViT/CNN-based architecture that implements CPRS by treating evolving CSI matrices as sequential image-like data, enabling efficient and adaptive transmission in dynamic environments.
-Simulation results using ray-tracing channel data generated in NVIDIA Sionna validate the proposed method, showing up to 36.60% throughput improvement over benchmark strategies.
-Index Terms:
-5G NR, Massive MIMO, Reference Signal, DM-RS, Channel Prediction, Deep Learning.
+We further propose a standards-compliant ViViT/CNN-based architecture that implements CPRS by treating evolving CSI matrices as sequential image-like data.
+This design enables efficient and adaptive transmission in dynamic environments.
+Ray-tracing-based simulations in NVIDIA Sionna validate the proposed method, demonstrating up to 36.60% throughput improvement over benchmark strategies.
 I Introduction
-The multiple-input-multiple-output (MIMO) systems [GESBERT2002] have been at the core of innovation since 4G LTE to enhance data throughput.
-However, despite its original goal, the introduction of massive MIMO has led to the challenge of increased channel dimensionality [LU2014].
+Multiple-input multiple-output (MIMO) systems [GESBERT2002] have been at the core of wireless innovation since 4G LTE.
+However, the introduction of massive MIMO has led to the challenge of increased channel dimensionality [LU2014].
 As the dimension of the channel matrix increases, both the transmission of the reference signal for channel estimation and the feedback of downlink channel state information (CSI) to the base station have become significant sources of overhead.
-The downlink CSI is critical for enhancing communication tasks such as beamforming [JAFAR2001], handover [TOLLI2008], and reference signal allocation [CAI2005].
+The downlink CSI is critical for enhancing communication operations such as beamforming [JAFAR2001], handover [TOLLI2008], and reference signal allocation [CAI2005].
 In time division duplexing (TDD) systems, the downlink channel can be inferred from the uplink channel due to channel reciprocity.
@@
 Within each slot, the PDSCH DM-RS can be configured to occupy 1 to 4 out of the 14 OFDM symbols.
-Recently, deep learning has emerged as a breakthrough solution for reducing CSI feedback overhead and is also being discussed in the 5G Advanced standards following 3GPP Rel-18 [RP213599].
+Recently, deep learning has emerged as a promising solution for reducing CSI feedback overhead and has been actively discussed for incorporation into 5G-Advanced standards following 3GPP Rel-18 [RP213599].
 The solutions can be broadly classified into two categories: CSI compression [WEN2018] and channel prediction [WANGJIE2019].
-We focus on channel prediction, which estimates downlink CSI from uplink CSI or data signals, even without explicit CSI feedback.
-The network architectures have evolved from CNNs [WANGJIE2019] to LSTMs [HELMY2023], and more recently, transformers [ZHOU2024].
-There have also been attempts to jointly optimize beamforming and handover, the primary objectives of channel prediction.
+We focus on channel prediction, which estimates downlink CSI from uplink CSI or data signals, even without explicit feedback.
+The network architectures have evolved from convolutional neural networks (CNNs) [WANGJIE2019] to long short-term memory (LSTM) models [HELMY2023], and more recently, transformers [ZHOU2024].
+There have also been attempts to jointly optimize beamforming and handover.
 For instance, [CHU2022] proposed a deep reinforcement learning algorithm for channel prediction and beamforming, while [JIANG2022] used a transformer-based approach for beamforming with predicted CSI.
-In terms of handover, channel quality prediction has been used to guide handover decisions [SKABA2024].
-However, to the best of the authors’ knowledge, no prior study has explored the critical issue of reference signal allocation in conjunction with channel prediction, despite their strong interdependence.
+Channel quality prediction has also been used to guide handover decisions [SKABA2024].
+However, to the best of the authors’ knowledge, no prior study has addressed the critical issue of reference signal allocation in conjunction with channel prediction, despite their strong interdependence.
 We address the allocation of the physical downlink shared channel (PDSCH) demodulation reference signal (DM-RS) within the NR frame structure, as illustrated in Fig.
@@
 Moreover, unlike other reference signals transmitted over the control plane, the DM-RS is carried on the data plane alongside data symbols.
-The NR standard provides various configuration options for the flexible placement of 1 to 4 DM-RS symbols within each slot, the transmission time interval (TTI) of the NR frame, as detailed in Table I, offering a level of flexibility not available in previous LTE standards.
-This design accommodates practical challenges such as mobility-induced channel aging, noise, and estimation uncertainty.
-The in-slot placement makes DM-RS allocation particularly important, as increasing the number of DM-RS symbols improves channel estimation accuracy at the expense of reducing transmittable data symbols, highlighting an inherent trade-off in the allocation design.
+The NR specification supports 1–4 DM-RS symbols per slot with multiple time-domain patterns (see Table I), offering flexibility to accommodate mobility, noise, and channel estimation uncertainty.
+The in-slot placement makes DM-RS allocation particularly important.
+Increasing the number of DM-RS symbols improves channel estimation accuracy but reduces the number of transmittable data symbols, highlighting an inherent trade-off in the allocation design.
 Therefore, the DM-RS must be optimally allocated to balance channel estimation quality and the number of available data symbols, considering the current downlink CSI.
@@
 DM-RS Length DM-RS positions pos0 pos1 pos2 pos3 Single-Symbol $l_{0}$ $l_{0}$, $l_{1}$ $l_{0}$, 7, 11 $l_{0}$, 5, 8, 11 Double-Symbol $l_{0}$ $l_{0}$, 10 - -
-This paper proposes a channel prediction-based reference signal allocation method, with a focus on DM-RS, for massive MIMO systems in B5G networks.
-While designed with B5G scenarios in mind, the proposed method remains fully compatible with the 5G NR frame structure and is applicable under current standard specifications.
-Contribution: 1) This paper introduces the concept of channel prediction-based reference signal allocation (CPRS), with a particular focus on DM-RS, as a practical approach to reducing CSI feedback overhead in massive MIMO FDD scenarios.
-2) We propose a ViViT/CNN-based [ARNAB2021] CPRS algorithm, designed with consideration of the NR frame structure, which interprets time-varying uplink channel matrices as video data.
+This paper proposes a joint optimization framework for channel prediction and reference signal allocation, with a particular focus on DM-RS, instead of treating them as independent modules.
+Although motivated by B5G scenarios, the proposed approach is fully compatible with the 5G NR frame structure and applicable under current standard specifications.
+Contribution: 1) This paper introduces the concept of channel prediction-based reference signal allocation (CPRS), focusing on DM-RS as a practical means of reducing CSI feedback overhead in massive MIMO FDD scenarios.
+2) We propose a video vision transformer (ViViT)/CNN-based [ARNAB2021] CPRS algorithm, designed with consideration of the NR frame structure, which interprets time-varying uplink channel matrices as video data.
 3) We generate ray-tracing-based channel data and perform simulations using NVIDIA Sionna [NVIDIA], demonstrating the effectiveness of the proposed method.
 II System Model and Problem Formulation
-We consider a massive MIMO-OFDM communication system, focusing on a downlink scenario with UE mobility, where the BS and UE communicate based on the NR frame structure.
+We consider a massive MIMO orthogonal frequency division multiplexing (OFDM) communication system, focusing on a downlink scenario with UE mobility, where the BS and UE communicate based on the NR frame structure.
 The numbers of receive antennas, transmit antennas, and subcarriers are denoted by $N_{\textnormal{R}}$, $N_{\textnormal{T}}$, and $N_{\textnormal{C}}$, respectively.
-The antennas at the UE receive the same transmitted signal, and the received signal at the UE for the $t$-th OFDM symbol in the slot and $k$-th subcarrier can be expressed as follows:
+The antennas at the UE receive the same transmitted signal, and the received signal for the $t$-th OFDM symbol in the slot and $k$-th subcarrier can be expressed as follows:
 $y_{t,k}=\mathbf{h}_{\textnormal{DL},t,k}\,\mathbf{v}_{t,k}\,x_{t,k}+\mathbf{w}_{t,k},$ (1)
 where $\mathbf{h}_{\textnormal{DL},t,k}\in\mathbb{C}^{N_{\textnormal{R}}\times N_{\textnormal{T}}}$ is the downlink channel matrix associated with the $t$-th symbol and $k$-th subcarrier, $\mathbf{v}_{t,k}\in\mathbb{C}^{N_{\textnormal{T}}\times 1}$ is the precoding vector, $x_{t,k}\in\mathbb{C}$ is the transmit symbol, and $\mathbf{w}_{t,k}\in\mathbb{C}^{N_{\textnormal{R}}\times 1}$ denotes additive noise.
-Then, each $t$-th symbol within the slot has a distinct downlink CSI as follows:
-$\mathbf{H}_{\textnormal{DL},t}=\left[\mathbf{h}_{\textnormal{DL},t,1}\,\mathbf{h}_{\textnormal{DL},t,2}\,\cdots\,\mathbf{h}_{\textnormal{DL},t,N_{{\textnormal{C}}}}\right]\in\mathbb{C}^{N_{\textnormal{R}}\times N_{\textnormal{T}}\times N_{\textnormal{C}}}$ (2)
-We define the overall downlink CSI for a slot as $\mathbf{H}_{\textnormal{DL}}$, which consists of the CSI matrices corresponding to all 14 OFDM symbols in the slot. It is expressed as:
-$\mathbf{H}_{\textnormal{DL}}=\left\{\mathbf{H}_{\textnormal{DL},t}\right\}_{t=1}^{14}$ (3)
+Each $t$-th OFDM symbol has a corresponding downlink CSI matrix,
+$\mathbf{H}_{\textnormal{DL},t}=\left[\mathbf{h}_{\textnormal{DL},t,1}\,\mathbf{h}_{\textnormal{DL},t,2}\,\cdots\,\mathbf{h}_{\textnormal{DL},t,N_{{\textnormal{C}}}}\right]\in\mathbb{C}^{N_{\textnormal{R}}\times N_{\textnormal{T}}\times N_{\textnormal{C}}}.$ (2)
+We define the collection of downlink CSI across the 14 OFDM symbols in a slot as
+$\mathbf{H}_{\textnormal{DL}}=\left\{\mathbf{H}_{\textnormal{DL},t}\right\}_{t=1}^{14}.$ (3)
 In our system, the BS is assumed to predict $\mathbf{H}_{\textnormal{DL}}$ using the uplink CSI from the current slot, denoted as $\mathbf{H}_{\textnormal{UL}}=\{\mathbf{H}_{\textnormal{UL},i}\in\mathbb{C}^{N_{\textnormal{T}}\times N_{\textnormal{R}}\times N_{\textnormal{C}}}\}_{i\in I_{u}}$, where $I_{u}$ is the set of DM-RS symbol indices in the current uplink slot.
 We assume that the uplink DM-RS is transmitted at four symbols per slot and $I_{u}=\{2,5,8,11\}$.
-The downlink channel prediction using a predictor function $f(\cdot)$ is formulated as:
-$\hat{\mathbf{H}}_{\textnormal{DL}}=f(\mathbf{H}_{\textnormal{UL}})$ (4)
+The downlink channel prediction using a predictor function $f(\cdot)$ is formulated as
+$\hat{\mathbf{H}}_{\textnormal{DL}}=f(\mathbf{H}_{\textnormal{UL}}).$ (4)
 The BS utilizes $\hat{\mathbf{H}}_{\textnormal{DL}}$ to identify the optimal allocation $p^{*}\in\mathcal{P}$, where $\mathcal{P}$ denotes the set of possible allocation configurations from which the BS selects one to configure the next slot, in order to maximize data throughput.
@@
 The objective function of (5) represents the data throughput (bits per slot), following the model in [DAS2015].
-We focus on the core optimization problem, excluding retransmissions as they fall outside the scope of this study.
-Here, $R$, $\textnormal{D}_{\textnormal{sent}}$, and $n_{s}$ are constants for a fixed numerology and thus independent of $p$, while $n_{p}$, the number of DM-RS symbols, is directly determined by the allocation $p$.
-The key challenge lies in modeling the complex relationship among $p$, $\mathbf{H}{\textnormal{UL}}$, and the resulting BLER at the UE in the next slot.
+For a fixed numerology, $R$, $\textnormal{D}_{\textnormal{sent}}$, and $n_{s}$ are constants, while $n_{p}$ is determined by the allocation $p$.
+Thus, the main difficulty is to model the relationship among $p$, $\mathbf{H}_{\textnormal{UL}}$, and the resulting BLER at the UE in the next slot.
 Even if the BS accurately estimates the $\mathbf{H}_{\textnormal{DL}}$, it remains challenging to determine the extent of BLER caused by channel errors in symbols between DM-RS placements.
 We adopt a data-driven approach to determine the relationship among $p$, $\mathbf{H}_{\textnormal{UL}}$, BLER, and ultimately, data throughput.
-Utilizing neural network algorithms often applied in nonconvex optimization, we propose an end-to-end method that jointly optimizes both the channel prediction problem in (4) and the DM-RS allocation in (5), classifying optimal allocation $p^{*}\in\mathcal{P}$ directly from $\mathbf{H}_{\textnormal{UL}}$.
-For the neural network-based classifier $\mathcal{F}(\cdot;\Theta)$ and estimated allocation $\hat{p}^{*}\in\mathcal{P}$, the proposed CPRS method is defined as:
-$\hat{p}^{*}=\mathcal{F}(\mathbf{H}_{\textnormal{UL}};\Theta)$ (6)
+Leveraging neural network algorithms commonly used for nonconvex optimization, we propose an end-to-end method that jointly optimizes the channel prediction problem in (4) and the DM-RS allocation in (5).
+In particular, the proposed CPRS framework directly classifies the optimal allocation $p^{*}\in\mathcal{P}$ from $\mathbf{H}_{\textnormal{UL}}$.
+For the neural network-based classifier $\mathcal{F}(\cdot;\Theta)$ and estimated allocation $\hat{p}^{*}\in\mathcal{P}$, the proposed CPRS method is defined as
+$\hat{p}^{*}=\mathcal{F}(\mathbf{H}_{\textnormal{UL}};\Theta).$ (6)
 We assume the possible allocation set $\mathcal{P}$ follows the NR standard in Table I.
@@
 III ViViT/CNN-based CPRS Network
-Fig.
-2 illustrates our proposed ViViT-based CPRS network, designed to predict optimal DM-RS allocations by exploiting the temporal evolution of uplink channel matrices $\mathbf{H}_{\textbf{UL}}$.
-We also introduce a lower-complexity CNN-based variant with the same input–output configuration, whose standard architecture is omitted for brevity.
-Each $\mathbf{H}_{\textbf{UL},i}$ for $i\in\{2,5,8,11\}$ is first decomposed into real and imaginary parts, which are concatenated to form a multi-channel representation, then reshaped into non-overlapping patches that are flattened and embedded as tokens.
-Specifically, the channel matrix, originally of size $N_{\textnormal{T}}\times N_{\textnormal{R}}\times N_{\textnormal{C}}$, is reshaped by flattening the antenna dimensions, resulting in the following input tensor $\mathbf{X}_{i}$ for the network:
-$\mathbf{H}_{\textbf{UL},i}\in\mathbb{C}^{N_{\textnormal{T}}\times N_{\textnormal{R}}\times N_{\textnormal{C}}}\quad\Longrightarrow\quad\mathbf{X}_{i}\in\mathbb{R}^{N_{\textnormal{T}}N_{\textnormal{R}}\times N_{\textnormal{C}}\times 2}$ (7)
-This transformation enables the network to process spatial features jointly along the flattened antenna dimension, while preserving the temporal structure along the subcarrier axis $N_{\textnormal{C}}$.
-Moreover, it maintains compatibility with standard convolutional operations without losing the inherent spatial correlations present in the original matrix.
-We then partition $\mathbf{X}_{i}$ into a set of non-overlapping patches and flatten each patch to form the initial tokens:
-$\mathbf{z}_{i}^{(0)}=\bigl{\{}\mathbf{x}_{i,1},\,\mathbf{x}_{i,2},\,\dots,\mathbf{x}_{i,P}\bigr{\}},$ (8)
-where $\mathbf{x}_{i,p}$ denotes the $p$-th patch and $P$ is the total number of patches. Learnable positional encodings are applied to the patch embeddings to incorporate spatial order.
+The CPRS framework predicts the optimal DM-RS allocation by exploiting the temporal evolution of the uplink channel matrices $\mathbf{H}_{\textnormal{UL}}$ and is inherently model-agnostic, allowing the use of various backbone architectures.
+In this paper, we instantiate CPRS using two empirically strong backbones: ViViT and a CNN model.
+ViViT provides global spatio–temporal modeling and preserves channel structures that are critical for DM-RS allocation.
+The CNN-based variant maintains the same input–output structure while offering lower computational complexity; its standard architecture is omitted for brevity.
+Before being processed by the backbone model, each uplink channel matrix $\mathbf{H}_{\textnormal{UL},i}$ for $i\in\{2,5,8,11\}$ is decomposed into real and imaginary parts and reshaped into a unified real-valued input tensor:
+$\mathbf{H}_{\textnormal{UL},i}\in\mathbb{C}^{N_{\textnormal{T}}\times N_{\textnormal{R}}\times N_{\textnormal{C}}}\;\Longrightarrow\;\mathbf{X}_{i}\in\mathbb{R}^{N_{\textnormal{T}}N_{\textnormal{R}}\times N_{\textnormal{C}}\times 2},$ (7)
+which preserves both spatial ($N_{\textnormal{T}}N_{\textnormal{R}}$) and subcarrier-domain ($N_{\textnormal{C}}$) structure. The four channel tensors are then stacked as
+$\mathbf{X}=\bigl[\mathbf{X}_{2},\mathbf{X}_{5},\mathbf{X}_{8},\mathbf{X}_{11}\bigr]\in\mathbb{R}^{T_{u}\times N_{\textnormal{T}}N_{\textnormal{R}}\times N_{\textnormal{C}}\times 2},$ (8)
+where $T_{u}=4$ denotes the number of UL DM-RS symbols per slot. Each $\mathbf{X}_{i}$ is partitioned into non-overlapping patches of size $(h_{p},w_{p})$, yielding
+$P=\frac{N_{\textnormal{T}}N_{\textnormal{R}}}{h_{p}}\cdot\frac{N_{\textnormal{C}}}{w_{p}},$ (9)
+where $P$ is the total number of patches. Let $\mathbf{X}_{i,p}\in\mathbb{R}^{h_{p}\times w_{p}\times 2}$ denote the $p$-th patch at time index $i$. Each patch is flattened and embedded as
+$\mathbf{e}_{i,p}=\mathbf{W}_{\mathrm{emb}}\,\mathrm{vec}(\mathbf{X}_{i,p})+\mathbf{b}_{\mathrm{emb}}+\mathbf{p}_{p},$ (10)
+where $\mathbf{W}_{\mathrm{emb}}$ and $\mathbf{b}_{\mathrm{emb}}$ are learnable projection parameters and $\mathbf{p}_{p}$ is a positional encoding. The resulting token matrix for time index $i$ is
+$\mathbf{Z}_{i}^{(0)}=\bigl[\mathbf{e}_{i,1}^{\top},\dots,\mathbf{e}_{i,P}^{\top}\bigr]^{\top}\in\mathbb{R}^{P\times D},$ (11)
+where $D$ denotes the token embedding dimension.
 Spatial Transformer Encoder
-In the spatial transformer stage, each set of token embeddings $\mathbf{z}_{i}^{(0)}$ for a given time index $i$ passes through a standard transformer encoder block.
-Concretely, let $\mathbf{Z}_{i}^{(\ell)}\in\mathbb{R}^{P\times D}$ be the output of the $\ell$-th encoder layer, where $D$ is the embedding dimension of each token.
-Each encoder layer consists of multi-head self-attention (MHSA) and a position-wise feedforward network (FFN), each followed by layer normalization.
-The MHSA operation is defined by:
-$\text{Attention}(Q,K,V)=\text{softmax}\!\Bigl{(}\frac{QK^{\top}}{\sqrt{d}}\Bigr{)}\!V,$ (9)
-where $Q$, $K$, and $V$ are linear projections of the input $\mathbf{Z}_{i}^{(\ell-1)}$, and $d$ is the dimensionality of each head.
-Multiple attention heads are concatenated to capture diverse representation subspaces.
-After processing all four channel matrices in a slot independently in the spatial transformer, we obtain spatially enhanced tokens for each time $i$.
+Each token matrix $\mathbf{Z}_{i}^{(0)}\in\mathbb{R}^{P\times D}$ is processed by $L_{s}$ spatial transformer layers.
+Each layer consists of multi-head self-attention (MHSA) and a feedforward network (FFN).
+The attention operation is
+$\mathrm{Attention}(Q,K,V)=\mathrm{softmax}\!\Bigl(\tfrac{QK^{\top}}{\sqrt{d}}\Bigr)V,$ (12)
+where $Q,K,V$ are linear projections of the input and $d$ is the per-head dimension. The layer updates are
+$\widetilde{\mathbf{Z}}_{i}^{(\ell)}$ $=\mathbf{Z}_{i}^{(\ell-1)}+\mathrm{MHSA}\!\bigl(\mathrm{LN}(\mathbf{Z}_{i}^{(\ell-1)})\bigr),$ (13)
+$\mathbf{Z}_{i}^{(\ell)}$ $=\widetilde{\mathbf{Z}}_{i}^{(\ell)}+\mathrm{FFN}\!\bigl(\mathrm{LN}(\widetilde{\mathbf{Z}}_{i}^{(\ell)})\bigr),$ (14)
+where $\ell=1,\dots,L_{s}$ and $\mathrm{LN}(\cdot)$ is layer normalization.
 Temporal Transformer Encoder
-We then stack the spatially transformed tokens across four time indices and feed them into the temporal transformer encoder.
-The same MHSA and FFN blocks now operate along the temporal dimension to learn correlations across DM-RS time positions.
-Stacking multiple temporal attention layers allows the network to capture channel evolution over time slots.
-After the temporal transformer encoder, the final token representations are pooled and passed through a multi-layer perceptron (MLP) head to predict the optimal downlink DM-RS allocation index.
-This prediction leverages the spatio-temporal correlations learned via the transformer’s self-attention mechanism, which effectively captures long-range dependencies, to complete the end-to-end decision pipeline.
-The entire ViViT/CNN-based CPRS network is trained using a cross-entropy loss, aligning the model’s predictions with the ground-truth optimal DM-RS allocation labels.
-The total computational complexity of the proposed ViViT model is $\mathcal{O}(L\cdot(N^{2}D+ND^{2}))$, where $L$ denotes the number of transformer layers, $N$ is the number of input tokens, and $D$ is the embedding dimension.
-For comparison, the CNN-based baseline consists of three 3D convolutional layers, each with computational complexity $\mathcal{O}(F\cdot C\cdot K^{3}\cdot H\cdot W\cdot T)$, where $K$ is the kernel size, $C$ the number of input channels, $F$ the number of output filters, and $H\times W\times T$ the spatial-temporal dimensions.
-Based on FLOP-level complexity analysis and efficient quantization techniques such as [ZMORA2021], both models are expected to be deployable at the microsecond level in a slot-based real-time system with parallel implementation.
-[^1] With $L{=}8$, $N{=}64$, and $D{=}192$ as in our experiments, the ViViT model requires $\sim\!239.2\times 10^{6}$ FLOPs.
-Assuming INT8 quantization on an NVIDIA A100 (624 TFLOPs), the theoretical inference time is $\sim\!3.83~{}\mu\text{s}$.
-The CNN model, with fewer FLOPs, can run even faster.
+The spatially processed tokens $\mathbf{Z}_{i}^{(L_{s})}$ are averaged across patches to obtain a $D$-dimensional summary:
+$\mathbf{u}_{i}=\tfrac{1}{P}\sum_{p=1}^{P}\mathbf{Z}_{i,p}^{(L_{s})}\in\mathbb{R}^{D},$ (15)
+where $\mathbf{Z}_{i,p}^{(L_{s})}$ is the $p$-th row (patch token) of $\mathbf{Z}_{i}^{(L_{s})}$. Stacking the four summaries yields the temporal sequence
+$\mathbf{U}^{(0)}=\bigl[\mathbf{u}_{2}^{\top},\mathbf{u}_{5}^{\top},\mathbf{u}_{8}^{\top},\mathbf{u}_{11}^{\top}\bigr]^{\top}\in\mathbb{R}^{T_{u}\times D}.$ (16)
+Temporal MHSA and FFN blocks are applied for $L_{t}$ layers:
+$\widetilde{\mathbf{U}}^{(\ell)}$ $=\mathbf{U}^{(\ell-1)}+\mathrm{MHSA}_{t}\!\bigl(\mathrm{LN}(\mathbf{U}^{(\ell-1)})\bigr),$ (17)
+[…]
+where $\mathrm{MHSA}_{t}$ and $\mathrm{FFN}_{t}$ operate along the temporal axis.
+A slot-level representation is obtained by temporal averaging:
+$\mathbf{h}=\tfrac{1}{T_{u}}\sum_{i\in\{2,5,8,11\}}\mathbf{U}_{i}^{(L_{t})}\in\mathbb{R}^{D},$ (19)
+and processed by an MLP:
+$\mathbf{o}$ $=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{h}+\mathbf{b}_{1})+\mathbf{b}_{2},$ (20)
+$\boldsymbol{\pi}$ $=\mathrm{softmax}(\mathbf{o}),$ (21)
+where $\boldsymbol{\pi}\in\mathbb{R}^{|\mathcal{P}|}$ is the predicted distribution over allocation patterns. The final allocation is
+$\hat{p}^{*}=\arg\max_{p\in\mathcal{P}}\pi_{p}.$ (22)
+[…]
 III-A Dataset Generation
@@
 After constructing a dynamic scene that reflects movement as illustrated in (a), perform ray tracing by positioning the BS and UE as shown in (b).
-We utilized the NVIDIA Sionna Ray Tracing to generate channel data [NVIDIA], as it more accurately reflects real-world propagation characteristics and physical environments than statistical channel models.
-Specifically, we employed the simple_street_canyon_with_cars scene included in the framework.
-As illustrated in Fig.
-3, this scenario features a road between buildings, where vehicles move along both lanes.
-All vehicles move at a uniform speed $v\in\{60,65,\dots,80\}$ km/h.
-The BS is installed on one of the buildings and the UE is positioned in one of the vehicles.
+We utilized the ray-tracing module of NVIDIA Sionna to generate channel data [NVIDIA], as it more accurately reflects real-world propagation characteristics and physical environments than statistical channel models.
+We employed the simple_street_canyon_with_cars scene, where vehicles move along a two-lane road between buildings, as shown in Fig.
+3.
+The BS is mounted on a building, and the UE is placed in a vehicle moving at $v\in\{60,65,\dots,80\}$ km/h.
 The BS has an 8×8 antenna array, while the UE has a 2×1 antenna array, forming a massive MIMO system.
@@
 The maximum number of rays was set to 20, with a maximum ray depth of 4 and a bandwidth of 100 MHz.
-For each velocity $v$, we obtained the uplink and downlink channel matrices for every symbol within 400 consecutive slots, each comprising 14 symbols and 64 subcarriers.
-Additionally, we augmented the original dataset by adding Gaussian noise with a magnitude of approximately 0.05 relative to the original channel matrices, in order to emulate practical channel estimation errors caused by uplink noise and other impairments.
-This augmentation generated an additional 4,000 slots.
-This resulted in a total dataset comprising 6,000 slots.
-Subsequently, for each slot, the optimal DM-RS allocation index was determined via exhaustive evaluation of all candidate configurations using Sionna link-level simulations, based on the objective function defined in (5).
-The dataset was then divided into training, validation, and test sets using an 8:1:1 ratio.
+To emulate practical estimation errors, we added Gaussian noise with magnitude $\approx$0.05 relative to the original channels, generating 4,000 additional slots and yielding 6,000 slots in total.
+For each slot, the optimal DM-RS allocation was obtained via exhaustive search using Sionna link-level simulations based on (5), and the dataset was split into training/validation/test sets with an 8:1:1 ratio.
+IV Experimental Results
+IV-A Considered Schemes
+CPRS (ViViT): The proposed scheme described in Section III.
+It adopts a tubelet embedding with a patch size of $(1,8,8)$ and a Transformer backbone consisting of 8 encoder layers, each with 16 attention heads and an embedding dimension of 128.
+These architectural parameters were selected based on ablation studies, where this configuration consistently achieved the best performance among the considered ViViT variants.
+CPRS (CNN): A CPRS variant based on a 3D convolutional neural network, where all $\mathbf{X}_{i}$ are concatenated into a four-channel input.
+The network consists of three Conv3D layers with 16, 32, and 64 filters, respectively, each followed by batch normalization and 3D max pooling, and a 128-unit dense layer at the output.
+This configuration was chosen as it yielded the highest performance among evaluated CNN-based architectures.
+Disjoint (ViViT/CNN): In the disjoint baseline, the downlink CSI is first predicted using a ViViT- or CNN-based channel predictor [YANG2020].
+To ensure a fair comparison, the ViViT predictor adopts the same patch size $(1,8,8)$ and backbone depth (8 layers with 16 attention heads) as CPRS (ViViT), while the CNN predictor uses the same 3D-CNN architecture as CPRS (CNN).
+The predicted CSI is then used to optimize the DM-RS pattern by maximizing channel estimation accuracy, following the procedure in [MASHHADI2021Prun] with adaptations for NR compliance.
+DRL [KIM2023]: A CNN-based deep Q-network employing an $\epsilon$-greedy policy ($\epsilon=0.3$), where the uplink CSI serves as the state, 3GPP-compliant DM-RS patterns constitute the action space, and throughput is used as the reward.
+Maximize Data Symbols: A heuristic scheme that maximizes the number of transmitted data symbols by minimizing the number of DM-RS symbols per slot ($n_{p}=1$).
+Random Average: A scheme that selects all possible DM-RS allocations with equal probability. The reported performance corresponds to the average over all possible allocations.
+Best: The ground-truth DM-RS allocation that achieves the optimal data throughput obtained through simulations.
 TABLE II: Classification accuracy comparison of different schemes for selecting the optimal DM-RS allocation index in the next slot, evaluated over 250 test slots.